OpenStudy (anonymous):

a/c = b/5 solve for c

OpenStudy (anonymous):

c=5a/b

OpenStudy (anonymous):

This was such an easy one, I know, but it has been FOREVER since I was in a math class and my brain just seemed to freeze up. Thank you. Can you show how to work it out. Again, I know it is simple, but I had a serious brain fart.

OpenStudy (anonymous):

a · 5=b · c b · c=a · 5 bc=a · 5 bc=5a c= 5a/b

OpenStudy (anonymous):

Ha! Thank you so much!

OpenStudy (anonymous):

np

OpenStudy (anonymous):

:)